OpenStudy (anonymous):

ill give a medal to the first person who can tell me how do u get the vertex of a parabola. be very easy in your anwer please. my problem is y=x^2+2x but i just want to know in general.

OpenStudy (anonymous):

ax^2+bx+c vertex(-b/(2a),f(-b/(2a)) -b/(2a)=-2/2=-1 f(-1)=(-1)^2+2(-1)=1-2=-1 vertex (-1,-1)

OpenStudy (anonymous):

do you know how to put a quadratic in standard form? completing the square?